OpenStudy (anonymous):

What is the slope of the line that passes through (c, d) and (c, 1/d)?

myininaya (myininaya):

Do you know the slope formula?

OpenStudy (anonymous):

yes. i got \[1\div d - d \div c-c \]

myininaya (myininaya):

Right but what is the bottom equal to?

OpenStudy (anonymous):

dont the two c's cross out?

OpenStudy (anonymous):

cuz yhu got c-c on the bottom

myininaya (myininaya):

What is c-c?

OpenStudy (anonymous):

the two points are (c,d) and (c, 1/d) so you have c-c

myininaya (myininaya):

which equals ?

myininaya (myininaya):

5-5=?

myininaya (myininaya):

6-6=? so what is c-c=?

OpenStudy (anonymous):

there is no numbers.. its LETTERS.. you suppose to solve

OpenStudy (anonymous):

the letters

myininaya (myininaya):

the c and d are numbers not letters

myininaya (myininaya):

If you have c of something and I take away that c of something, then how much of that something do you have left?

OpenStudy (anonymous):

\[m=\frac{ (y2-y1) }{ (x2-x1) }\]

OpenStudy (anonymous):

thanks :)

myininaya (myininaya):

Great. You got it?